OpenStudy (anonymous):

factors of 3528 that add up to 23?

myininaya (myininaya):

so we have xy=3528 and x+y=23. solve the sum for either x or y

myininaya (myininaya):

if you solve for y: y=23-x, so xy=x(23-x)=3528

myininaya (myininaya):

but now we have -x^2+23x=3528 so we didn't get anywhere because we still have to come up with factors of 3528 that add up to be 23. Do you know the quadratic formula?

OpenStudy (anonymous):

its a trinomial. i have to factor it, but i couldnt find any thing to pull out. heres the equation: \[63a^2+23ab-56b^2\]

myininaya (myininaya):

oh factors of -3528 that add up to be 23 one sec

myininaya (myininaya):

ok so i normally make a little tree and move numbers around and I got 72(-49)=-3528 and 72+(-49)=23

OpenStudy (anonymous):

ohhhh okay i see what i did wrong. having the negative right?

myininaya (myininaya):

right

OpenStudy (anonymous):

okayy i see, thank you!

myininaya (myininaya):

np

OpenStudy (anonymous):

so when i do it i write it as...

OpenStudy (anonymous):

\[63a^2+79ab-41ab-5b^2\]

myininaya (myininaya):

63a^2+72ab-49ab-5b^2 or 63a^2-49ab+72ab-5b^2

OpenStudy (anonymous):

okay cool, thanks!

myininaya (myininaya):

well that 5 is suppose to be 56 in both cases